# The Coding Limits of Robust Watermarking for Generative Models

### v1

@@ 3.2 Impossibility of High Tampering Rates
A central question for messageless secret-key codes is how robust they can be to tampering: if an adversary is allowed to modify a fraction of the codeword, can the code reliably detect tampering and also maintain soundness? It turns out that these two properties are fundamentally at odds if the tampering is sufficiently strong.
To see the tension, consider the following intuition. If the set of possible tampering functions includes all constant functions (that is, for every string $\hat{\gamma}$, there is a tampering function that always outputs $\hat{\gamma}$), then the decoder faces a contradiction: soundness requires that, for almost every $\hat{\gamma}$, the decoder outputs $\mathtt{invalid}$; yet tamper detection requires that, for any $\hat{\gamma}$ produced by tampering with a valid codeword, the decoder outputs $\mathtt{tampered}$. Since the decoder must assign a single label to each string, it cannot satisfy both properties for the same input $\hat{\gamma}$.
This intuition extends to the case where the adversary is limited to changing only a fraction $\alpha$ of the codeword, even if changes are made independently at random across positions. We now make this precise.
Let $\Sigma$} be an alphabet of size $q>1$, and let $n$ denote the codeword length. Define $\mathcal{F}_{\alpha n}^{\mathrm{ind}}$ as the family of functions $f$ where, for each input $\gamma\in{\Sigma}^{n}$, $f(\gamma)$ is produced by independently changing each symbol: for each position, the symbol is replaced with a new random symbol from $\Sigma$} with probability p=11/qp=1-1/q, and remains unchanged with probability $1/q$. The expected fraction of changed positions is exactly 11/q1-1/q.
@@
Let $\Gamma$ be any messageless secret-key code over $\Sigma$} with codeword length $n\geq 1$ and $|{\Sigma}|=q>1$. Suppose $\Gamma$ satisfies tamper detection for the family $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ with $\alpha=(1-1/q)(1+\delta)$ for some $\delta\in(0,1)$. Then $\Gamma$ cannot satisfy soundness.
@@
Let $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})$ and let $\gamma=\mathsf{Enc}(\mathit{sk})$ be an honestly generated codeword. Consider the tampering function $f$ that acts independently on each coordinate as follows: for each $i\in\{1,\ldots,n\}$, set
[…]
Let $X$ be the number of positions where $f(\gamma)$ differs from $\gamma$. The random variables $X_{i}=\mathbf{1}\{(f(\gamma))_{i}\neq\gamma_{i}\}$ are independent Bernoulli variables with $\mathbb{E}[X_{i}]=1-1/q$, so $\mathbb{E}[X]=n(1-1/q)$.
@@
which is negligible in $n$. Thus, with overwhelming probability, $f$ changes at most $\alpha n=(1-1/q)(1+\delta)n$ positions, so $f\in\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ except with negligible probability.
@@
Now, notice that $f(\gamma)$ is distributed exactly as a uniformly random string in ${\Sigma}^{n}$. This is because in each position, the symbol is resampled independently and uniformly, so the overall string is independent of $\gamma$.
@@
By the soundness property, for a randomly chosen $\hat{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}{\Sigma}^{n}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}}}\right]\leq\mathsf{negl}(\lambda)$.
@@
By tamper detection, for any honestly generated codeword $\gamma$ and any $f\in\mathcal{F}^{\mathrm{ind}}_{\alpha n}$, and for $\tilde{\gamma}=f(\gamma)$ with $\tilde{\gamma}\neq\gamma$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mathtt{tampered}}}\right]\leq\mathsf{negl}(\lambda)$. But since $f(\gamma)$ is distributed as $\hat{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}{\Sigma}^{n}$ and is almost always different from $\gamma$, we have for a randomly chosen $\hat{\gamma}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})=\mathtt{tampered}}}\right]\allowbreak\geq 1-\mathsf{negl}(\lambda)$.
Therefore, for most random strings $\hat{\gamma}$, the decoder must output both $\mathtt{invalid}$ (by soundness) and $\mathtt{tampered}$ (by tamper detection) except with negligible probability. This is impossible, since the decoder can only output one of these values for each input.
@@
We conclude that no messageless secret-key code can be sound and $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper-detecting at tampering rates $\alpha\geq(1-1/q)(1+\delta)$. ∎
This bound is tight for the family of independent symbol tampering functions, and matches intuition from coding theory: if too large a fraction of the codeword can be arbitrarily changed, no procedure can reliably distinguish random tampering from invalid codewords. In the case of binary codewords (that is, ${\Sigma}=\{0,1\}$), the impossibility threshold becomes particularly simple:
@@
Let $\Gamma$ be any messageless secret-key code with alphabet $\{0,1\}$ and codeword space $\{0,1\}^{n}$ for $n\geq 1$. If $\Gamma$ satisfies $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper detection for $\alpha=(1+\delta)/2$ for any $\delta\in(0,1)$, then $\Gamma$ cannot also satisfy soundness.
@@
These impossibility results do not depend on whether the decoder ever accepts any honestly generated codeword as valid. For instance, a code could be designed so that the decoder never outputs $\mathtt{valid}$ at all; it would trivially satisfy soundness (since random strings are always labeled invalid) and could still claim to detect tampering. Our theorems show that, even allowing for such “incorrect” codes, soundness and tamper detection cannot both be satisfied at high tampering rates. This is why correctness is not assumed or required for our lower bounds.
Remark 1.
It’s worth noting that the limitations described above are even more robust than they may first appear. First, suppose we relax our expectations for soundness, allowing the decoder to make mistakes on a constant fraction $\epsilon$ of random inputs, rather than insisting on negligible error. The underlying tension remains: once the adversary is able to tamper with a sufficiently large fraction of the codeword, the code can only be sound for a fraction $\epsilon$ of random strings if it tolerates a correspondingly large tamper-detection error—specifically, at least $(1-\epsilon)\big{(}1-\exp(-\frac{\delta^{2}n}{3}(1-1/q))\big{)}$ for tampering rates $\alpha=(1-1/q)(1+\delta)$ and any $\delta\in(0,1)$. In essence, even a “lenient” code, willing to make plenty of mistakes, still cannot robustly detect high-rate tampering.
Second, these limits do not depend on the codeword length being fixed in advance. If codewords are allowed to be any length, the adversary can simply determine the length and tamper independently with the appropriate fraction of positions, and the same contradiction arises. The impossibility persists no matter how the codeword length is chosen.
Taken together, these points highlight a fundamental barrier that goes beyond the classical bounds from error-correcting codes. Here, the impossibility is a direct consequence of trying to satisfy both soundness and strong tamper detection, even in the highly flexible messageless setting.
3.3 Supporting Messages
Just as in the messageless case, we can define secret-key codes that are designed not only to detect tampering, but also to carry and recover messages. Here, the code serves two roles at once: if a codeword has not been altered, it should decode back to the original message; if the codeword has been tampered with or is otherwise invalid, the decoder should indicate this fact.
Formally, such a code consists of three algorithms: a key generator $\mathsf{KGen}$, an encoder $\mathsf{Enc}$ that takes both a secret key and a message and outputs a codeword, and a decoder $\mathsf{Dec}$ that uses the secret key to recover either the original message, or a special symbol ($\mathtt{invalid}$ or $\mathtt{tampered}$) indicating something has gone wrong.
As before, we will be interested in the fundamental properties of such codes: correctness, soundness, tamper detection, and pseudorandomness. The presence of a message simply means that correctness now asks for exact message recovery from an unaltered codeword, and the other properties extend in a natural way. For completeness, we give their formal definitions below.
Formally, a secret-key multi-message code is defined by a triple of polynomial-time algorithms $\Gamma=(\mathsf{KGen},\mathsf{Enc},\mathsf{Dec})$, specified as follows:
$\mathsf{KGen}(1^{\lambda})$:
A randomized key-generation algorithm that, on input the security parameter, outputs a secret key $\mathit{sk}$.
$\mathsf{Enc}(\mathit{sk},\mu)$:
A (possibly randomized) encoding algorithm that, given the secret key $\mathit{sk}$ and a message $\mu\in{{\Sigma}^{m}}$, outputs a codeword $\gamma\in{\Sigma}^{n}$.
$\mathsf{Dec}(\mathit{sk},\gamma)$:
A deterministic decoding algorithm that, given the secret key $\mathit{sk}$ and a codeword $\gamma$, outputs either a message in ${\Sigma}^{m}$}, or one of the symbols $\mathtt{invalid}$ or $\mathtt{tampered}$.
The key properties for such codes are defined as follows:
Definition 9 (Correctness).
A secret-key multi-message code $\Gamma$ satisfies correctness if for all $\lambda\in\mathbb{N}$, all $\mathit{sk}\in\mathsf{KGen}(1^{\lambda})$, and all messages $\mu\in{{\Sigma}^{m}}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\mathsf{Enc}(\mathit{sk},\mu))=\mu}}\right]=1$, where the probability is taken over the randomness of the encoding, if any.
Definition 10 (Soundness).
A secret-key multi-message code $\Gamma$ satisfies soundness if for every fixed codeword $\hat{\gamma}\in{\Sigma}^{n}$, $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]\leq\mathsf{negl}(\lambda)$.
Definition 11 ($\mathcal{F}$-Tamper Detection).
$\Gamma$ satisfies $\mathcal{F}$-tamper detection if for any tampering function $f\in\mathcal{F}$ and any message $\mu\in{{\Sigma}^{m}}$,
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mathtt{tampered}\wedge\tilde{\gamma}\neq\gamma:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda}),\ \gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Enc}(\mathit{sk},\mu),\ \tilde{\gamma}=f(\gamma)}}\right]\leq\mathsf{negl}(\lambda).$
Definition 12 (Pseudorandomness).
$\Gamma$ satisfies pseudorandomness if for every PPT distinguisher $\mathsf{D}$, $\mathopen{}\mathclose{{\left|\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{D}^{\mathcal{O}_{\sf real}}(1^{\lambda})=1:\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})}}\right]-\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{D}^{\mathcal{O}_{\sf rand}}(1^{\lambda})=1}}\right]}}\right|\leq\mathsf{negl}(\lambda)$, where $\mathcal{O}_{\sf real}(\mu)$ returns $\mathsf{Enc}(\mathit{sk},\mu)$ and $\mathcal{O}_{\sf rand}(\mu)$ returns a uniformly random string from ${\Sigma}^{n}$.
Remark 2.
Our impossibility results extend from messageless codes to codes that also carry messages. The reason is simple: if you have a code that can both recover messages and detect tampering, you can always just ignore the message part and use the code to check whether an input is valid, tampered, or invalid. In other words, any code that solves the harder message-recovery problem automatically solves the simpler messageless problem as well. So if it is impossible to achieve soundness and tamper detection beyond a certain threshold without messages, it is also impossible when messages are present.
3.4 From Tamper Detection to Error Correction
The barrier we have identified for tamper detection in messageless codes also holds in the stronger setting of error correction for codes that encode messages. This is because error correction necessarily includes the task of detecting whether a codeword has been tampered with: if a code can recover the message, it can certainly tell when decoding fails.
Formally, a code achieves error correction for a family of tampering functions $\mathcal{F}$ if, whenever a codeword $\gamma$ encoding a message $\mu$ is tampered with to produce $\tilde{\gamma}\neq\gamma$, the decoder outputs the original message $\mu$ except with negligible probability:
Definition 13 ($\mathcal{F}$-Error Correction).
A secret-key multi-message code $\Gamma$ achieves $\mathcal{F}$-error correction if, for all $f\in\mathcal{F}$ and all $\mu$, when $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{KGen}(1^{\lambda})$, $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{Enc}(\mathit{sk},\mu)$, and $\tilde{\gamma}=f(\gamma)$ with $\tilde{\gamma}\neq\gamma$, we have $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mu}}\right]\leq\mathsf{negl}(\lambda)$.
A code with correctness, soundness, error correction, and pseudorandomness is a pseudorandom error-correcting code (PRC) [CG24]. In our model, the decoder may output a tampered flag, but this does not affect the fundamental limits.
The key point is that impossibility for messageless codes immediately implies impossibility for message-carrying codes. If a code can correct errors for even a single message, it can be viewed as a messageless code by fixing that message and treating every successful decoding as “valid.” The converse also holds. To obtain a multi-message code from a single-message code, one can apply the standard transformation of Christ and Gunn [CG24], which converts any single-message PRC into a multi-message one, preserving soundness and error correction.
Theorem 3 (Extension of Impossibility).
Let $\mathcal{F}$ be any family of tampering functions. If no messageless secret-key code achieves soundness and $\mathcal{F}$-tamper detection above a given tampering rate, then no secret-key multi-message code (including PRCs) achieves error correction for messages above that rate.

### v2

@@ 3.2 Impossibility of High Tampering Rates
A central question for zero-bit tamper-detection codes is how robust they can be to tampering. If an adversary is allowed to modify a fraction of the symbols in a codeword, can the decoder still reliably detect that tampering has occurred while maintaining soundness on arbitrary strings? In this subsection we show that these two requirements come into fundamental conflict once the tampering rate is high enough.
To build some intuition, consider an extreme case in which the tampering family contains all constant functions: for every fixed string $\hat{\gamma}\in{\Sigma}^{n}$, there is a function $f_{\hat{\gamma}}$ that ignores its input and always outputs $\hat{\gamma}$. Soundness says that, over the choice of the key, the decoder should almost always label $\hat{\gamma}$ as $\mathtt{invalid}$. On the other hand, if $\hat{\gamma}$ can arise as $f_{\hat{\gamma}}(\gamma)$ for an honestly generated codeword $\gamma$, tamper detection would like the decoder to label $\hat{\gamma}$ as $\mathtt{tampered}$. Since the decoder can assign only one label to each string, it cannot satisfy both requirements for the same $\hat{\gamma}$. This already suggests that strong tamper detection is hard to reconcile with soundness once we allow very powerful tampering.
Our main result shows that a similar tension appears even for a much weaker, randomized tampering model in which each symbol is modified independently. Let $\Sigma$} be an alphabet of size $q>1$, and let $n$ denote the codeword length. We focus on the following independent tampering rule: for each input $\gamma\in{\Sigma}^{n}$, the adversary produces $\tilde{\gamma}$ by choosing each symbol of $\tilde{\gamma}$ independently and uniformly from $\Sigma$}, conditional on $\gamma$. For each position, the new symbol equals the original one with probability $1/q$ and differs from it with probability:
$1-\frac{1}{q}.$
Thus the expected fraction of changed positions is 11/q1-1/q. Standard concentration bounds imply that, for any fixed $\delta\in(0,1)$, with overwhelming probability the Hamming distance between $\gamma$ and $\tilde{\gamma}$ is at most:
$\Bigl(1-\frac{1}{q}\Bigr)(1+\delta)\,n.$
We write $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ for a family of tampering functions that includes this independent symbol rule and whose outputs differ from their inputs in at most $\alpha n$ positions with high probability.
With this notation in place we can state the main impossibility theorem.
@@
Let $\Gamma$ be any zero-bit tamper-detection code over $\Sigma$} with codeword length $n\geq 1$ and $|{\Sigma}|=q>1$. Suppose $\Gamma$ satisfies $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper detection for some
$\alpha=\Bigl(1-\frac{1}{q}\Bigr)(1+\delta)$
with $\delta\in(0,1)$. Then $\Gamma$ cannot satisfy soundness.
@@
Let $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})$ and let $\gamma=\mathsf{Enc}(\mathit{sk})$ be an honestly generated codeword. Consider the tampering function $f_{\gamma^{\prime}}$ defined as $f_{\gamma^{\prime}}(\gamma)=\gamma^{\prime}$ for every $\gamma\in{\Sigma}^{n}$ where $\gamma^{\prime}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}{\Sigma}^{n}$. Let $X$ be the number of positions where $f_{\gamma^{\prime}}(\gamma)$ differs from $\gamma$ and $(f_{\gamma^{\prime}}(\gamma))_{i}$ the $i$-th symbol output by $f_{\gamma^{\prime}}$ on input $\gamma$. The random variables $X_{i}=\mathbf{1}\{(f_{\gamma^{\prime}}(\gamma))_{i}\neq\gamma_{i}\}$ are independent Bernoulli variables with $\mathbb{E}[X_{i}]=1-1/q$, so $\mathbb{E}[X]=n(1-1/q)$.
@@
which is negligible in $n$. Thus, when $\gamma^{\prime}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}{\Sigma}^{n}$, with overwhelming probability $f_{\gamma^{\prime}}$ changes at most $\alpha n=(1-1/q)(1+\delta)n$ positions, so $f_{\gamma^{\prime}}\in\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ except with negligible probability.
@@
Now, notice that $f_{\gamma^{\prime}}(\gamma)$ is distributed exactly as a uniformly random string in ${\Sigma}^{n}$. This is because $f_{\gamma^{\prime}}(\gamma)=\gamma^{\prime}$ for every $\gamma\in{\Sigma}^{n}$ and $\gamma^{\prime}$ is randomly sampled from ${\Sigma}^{n}$.
By the soundness property, for every fixed $\hat{\gamma}\in{\Sigma}^{n}$,
@@
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}}}\right]\leq\mathsf{negl}(\lambda),$
where the probability is over $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})$. In particular, this bound also holds when $\hat{\gamma}$ is chosen uniformly at random, so for a randomly chosen $\hat{\gamma}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}{\Sigma}^{n}$ we still have $\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\hat{\gamma})\neq\mathtt{invalid}}}\right]\leq\mathsf{negl}(\lambda)$.
Moreover, by tamper detection, for any honestly generated codeword $\gamma$ and any $f\in\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ (including $f_{\gamma^{\prime}}$), and for $\tilde{\gamma}=f(\gamma)$ with $\tilde{\gamma}\neq\gamma$, we have
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\tilde{\gamma})\neq\mathtt{tampered}}}\right]\leq\mathsf{negl}(\lambda),$
@@
where the probability is over $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})$, $\gamma{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{Enc}(\mathit{sk})$, and the randomness of $f$. Since $f_{\gamma^{\prime}}(\gamma)$ (which is in $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$ with overwhelming probability) is distributed as $\gamma^{\prime}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}{\Sigma}^{n}$, it is almost always (except with negligible probability) different from $\gamma$. In turn, this means that for a randomly chosen $\gamma^{\prime}=f_{\gamma^{\prime}}(\gamma)$,
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\gamma^{\prime})=\mathtt{tampered}}}\right]\geq 1-\mathsf{negl}(\lambda).$
Therefore, under the joint distribution of $(\mathit{sk},\gamma^{\prime})$ with $\mathit{sk}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{KGen}(1^{\lambda})$ and $\gamma^{\prime}{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}{\Sigma}^{n}$, we have both
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\gamma^{\prime})=\mathtt{invalid}}}\right]$ $\geq 1-\mathsf{negl}(\lambda),$
$\mathbb{P}\mathopen{}\mathclose{{\left[\mathsf{Dec}(\mathit{sk},\gamma^{\prime})=\mathtt{tampered}}}\right]$ $\geq 1-\mathsf{negl}(\lambda),$
even though these events are disjoint. This is impossible, since the decoder can only output one of these values for each input.
@@
We conclude that no zero-bit tamper-detection code can be sound and $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper-detecting at tampering rates $\alpha\geq(1-1/q)(1+\delta)$. ∎
In words, once we ask for tamper detection against independent symbol modifications that, with high probability, corrupt slightly more than a (11/q)(1-1/q) fraction of the positions, there is no way to keep the decoder sound on arbitrary fixed strings. The value
$\alpha^{*}=1-\frac{1}{q}$
is the critical corruption rate in this model: below it, robust tamper detection is achievable (as we show with a matching construction), while above it, soundness and tamper detection cannot coexist.
In the binary case the statement takes a particularly clean form.
@@
Let $\Gamma$ be any zero-bit tamper-detection code with alphabet $\{0,1\}$ and codeword space $\{0,1\}^{n}$ for $n\geq 1$. If $\Gamma$ satisfies $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper detection for
$\alpha=\frac{1+\delta}{2}$
for some $\delta\in(0,1)$, then $\Gamma$ cannot also satisfy soundness.
@@
Notice that these impossibility results do not rely on correctness: nowhere in the proof do we use the requirement that an honestly generated codeword must be accepted as $\mathtt{valid}$. Even a degenerate scheme that never outputs $\mathtt{valid}$ would trivially satisfy soundness (since all random strings are labeled $\mathtt{invalid}$) and might still claim to detect tampering. The theorem shows that, at high tampering rates, such a scheme also cannot satisfy the $\mathcal{F}^{\mathrm{ind}}_{\alpha n}$-tamper detection property. This is why correctness is not assumed in our lower bounds.
Extension to weak soundness and variable length
The same phenomenon persists even if we weaken soundness substantially. Recall the definition of constant-bounded soundness (7): for a parameter $c\in(0,1]$, the decoder is allowed to misclassify a fixed string as anything other than $\mathtt{invalid}$ with probability up to $c$ over the choice of the key. For the independent tampering model above and corruption rates
$\alpha=\Bigl(1-\frac{1}{q}\Bigr)(1+\delta),$
one can show that if a code has constant-bounded soundness with parameter $c$, then for large $n$ its tamper-detection error under such tampering must be at least
$(1-c)\Bigl(1-\exp\bigl(-\tfrac{\delta^{2}}{3}(1-1/q)\,n\bigr)\Bigr),$
which tends to 1c1-c as $n$ grows. In particular, even a “coin-flip” decoder with $c=1/2$ cannot hope to detect high-rate independent tampering with small error; beyond the threshold $\alpha^{*}$ it must fail with probability close to $1/2$ on tampered codewords.
Finally, the impossibility does not depend on fixing the blocklength in advance. Allowing the code to use different lengths $n$ on different inputs does not help: for each length $n$ we can define the corresponding independent tampering rule on ${\Sigma}^{n}$, and the same argument goes through. In this sense the barrier identified by 1 is intrinsic to the combination of soundness and strong tamper detection, rather than an artifact of a particular choice of codeword length.
4 Connecting Watermarking and Tamper-detection Codes
The fundamental limits we have established for zero-bit tamper-detection codes also govern the robustness of watermarking schemes for generative models. To make this connection precise, we show how any watermarking scheme that is sound and robust against tampering can be used to build a secret-key code with exactly the same security guarantees.
Theorem 2.
Assume there exists a watermarking scheme $\Pi$ for a class of generative models $\mathcal{M}=\{\mathsf{Model}:\Sigma^{*}\rightarrow\Sigma^{*}\}$ satisfying soundness and $^\hat{\mathcal{F}}$-robustness (for some family $\hat{\mathcal{F}}$). [^2] For the sake of clarity, we assume the prompt and output space of $\mathsf{Model}$ is defined over the same alphabet. The result generalizes to different alphabets as well. Then, there exists a zero-bit tamper-detection code $\Gamma$ with alphabet space $\Sigma$} and codeword space $\Sigma$$^{*}$ satisfying soundness and $^\hat{\mathcal{F}}$-tamper detection.